OpenStudy (anonymous):

triangleABC has vertices A(–2, 4), B(6, 0), and C(–4, 0). Is triangleABC a right triangle? triangleABC is a right triangle because a. triangleABC is a right triangle because a. triangleABC is a right triangle because a. triangleABC is not a right triangle.

OpenStudy (anonymous):

slope of AB = (0-4)/(6+2) =-1/2 slope of AC =(0-4)(-4+2)=2 since solpe of AB * slope of AC = -1/2 *2 =-1 hence triangle ABC is rt angled at A

OpenStudy (anonymous):

also check BC^2= 100 and AB^2 =64+16 =80 and AC^2=16+4=20 thus BC^2=AB^2+AC^2 (Pythogarean rule)